OpenStudy (anonymous):

solve by completing the square. 4x^2+4x+3=0

OpenStudy (anonymous):

4x^2+4x+3=0 you'll get a = 4 b = 4 c = 3 subsitute all of them (a, b, and c) to quadratic equations

OpenStudy (anonymous):

yes do that lolz

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a } \]

OpenStudy (anonymous):

what else can do

OpenStudy (anonymous):

Check your equations Is your equations 4x^2-4x+3=0 4x^2+4x-3=0 or 4x^2-4x-3=0 ????

OpenStudy (anonymous):

4x^2+4x+3=0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2 - 4*a*c} }{ 2a } = \frac{ -4 \pm \sqrt{4^2 - 4*4*3} }{ 2*4 } = \frac{ -4 \pm \sqrt{-32} }{ 8 }\]

OpenStudy (anonymous):

please help. solve by completing the square 4x^2+4x+3=0

OpenStudy (anonymous):

\[x = \frac{ -4 \pm \sqrt{-1}*\sqrt{32} }{ 8} = \frac{ -4 \pm i*\sqrt{32} }{ 8 } = \frac{ -4 \pm 4\sqrt{2}*i }{ 8 }\]

OpenStudy (anonymous):

so you will get \[x = \frac{ -4 + 4\sqrt{2}*i }{ 8 } =\frac{ -1 + \sqrt{2}*i }{ 2 } = -\frac{ 1 }{ 2 } + \frac{ 1 }{ 2 } \sqrt{2} *i\]

OpenStudy (anonymous):

and

OpenStudy (anonymous):

\[x = \frac{ -4 - 4\sqrt{2} *i}{ 8 } = \frac{ -1 - \sqrt{2}*i }{ 2} = -\frac{ 1 }{ 2 } - \frac{ 1 }{ 2 } \sqrt{2}*i\]

OpenStudy (anonymous):

i did it for the lulz i did it for the lulz i did it for the lul

OpenStudy (anonymous):

get it @eogbonna ??

OpenStudy (anonymous):

help with solve by completing the square 4x^2+4x+3=0

OpenStudy (anonymous):

help ,i am running out of time